OpenStudy (anonymous):

write an equation of a line in slope intercept form going through the points (- 2, 0) and (4, - 3). Show your work

OpenStudy (ivettef365):

first you need to find the slope

OpenStudy (anonymous):

idk how to do any of this v~v

OpenStudy (ivettef365):

use the slope form; y2 - y1 ------- x2 - x1

OpenStudy (ivettef365):

x1,y1 x2 y2 (-2,0) and (4, -3)

OpenStudy (anonymous):

so -3 -0 / 4- -2

OpenStudy (anonymous):

?

OpenStudy (mertsj):

yes

OpenStudy (mertsj):

-3/6=-1/2 = slope

OpenStudy (anonymous):

i dont understand... can you explain

OpenStudy (mertsj):

Now use point slope form to find the equation of the line:

OpenStudy (mertsj):

\[y-y _{1}=m(x-x _{1})\]

OpenStudy (mertsj):

You may use either point. I, personally, would choose (-2,0) for (x_1.y+1)

OpenStudy (mertsj):

*y_1

OpenStudy (anonymous):

i dont know point slope

OpenStudy (mertsj):

That's why I posted it for you.

OpenStudy (anonymous):

-_-